OpenStudy (anonymous):

Determine the solution set for z: z6 - 4z5 - 30z4 = 0 -3, 5 0,-3, 5 3, -5 0, 3, -5

OpenStudy (vishweshshrimali5):

Hi there!

OpenStudy (anonymous):

Hi can you help me?

OpenStudy (vishweshshrimali5):

Yep! Ready?

OpenStudy (anonymous):

Yes

OpenStudy (vishweshshrimali5):

Okay! See, there are several methods of finding the solution set. Let's start with the trickiest one: Put the value of z from each option into the equation and see if the solution is valid. If yes, then you have your solution set, else continue to next option

OpenStudy (vishweshshrimali5):

Need an example ?

OpenStudy (anonymous):

Is the answer 0, -3, and 5

OpenStudy (vishweshshrimali5):

Give me a moment. I will check it ou

OpenStudy (anonymous):

Okay

OpenStudy (vishweshshrimali5):

z = 0 is correct. But, check again for the rest ones

OpenStudy (anonymous):

So it's the last choice?

OpenStudy (anonymous):

Am I right?

OpenStudy (vishweshshrimali5):

Are you sure the question is correct ? I mean can you please recheck the equation once ?

OpenStudy (anonymous):

Determine the solution set for z: z6 - 4z5 - 30z4 = 0 -3, 5 0,-3, 5 3, -5 0, 3, -5

OpenStudy (anonymous):

That's the question given to us

OpenStudy (vishweshshrimali5):

I think there is some mistake in question: See, the reason is this: \[\large{z^6 - 4z^5 - 30z^4 = 0}\] \[\large{z^4(z^2 - 4z - 30) = 0}\] Now, there are two possibilities for this: 1. \(\large{z^4 = 0} \implies z = 0\) 2. \(z^2 - 4z - 30 = 0 \implies z = 2 - \sqrt{34} \text{ or } z = 2 + \sqrt{34}\)

OpenStudy (vishweshshrimali5):

So, there is some mistake in the question or in the options

OpenStudy (anonymous):

I'm getting graded on this questions what do I choose?

OpenStudy (vishweshshrimali5):

I tell you what: Write "none of the options is correct" and go to your teacher and request for clarification. And, if possible, include the solution I wrote above. It will help you :)

OpenStudy (anonymous):

There's no option for it only pick a choice

OpenStudy (anonymous):

Can anyone else solve it and do they get the same as you?

OpenStudy (vishweshshrimali5):

Well, wolframalpha solved it and got the same thing... I would say go for (b) or (d) --> whichever one you want as both are wrong in reality :D

OpenStudy (anonymous):

I don't know what to do?!

Nnesha (nnesha):

plug solution into the equation if both sides are equal then those two number would be your answer let's see if its work...

OpenStudy (anonymous):

Idk how to!

OpenStudy (anonymous):

Is it any to the choices?

OpenStudy (anonymous):

I mean any of the choices?

Nnesha (nnesha):

like vishweshshrimali5 says . seems like there is an error for sure 0 is one of solution

Nnesha (nnesha):

the zeros he/she got are 2 -sqrt{34} and 2+sqrt{34} use calculator and change it to decimal you will get another hint :-)

OpenStudy (anonymous):

I got 7.83095189485

Nnesha (nnesha):

what about the first one ?

OpenStudy (anonymous):

-3.83095189485

OpenStudy (anonymous):

What do I do with those numbers?

Nnesha (nnesha):

2nd one is negative 3.83 seems like it can be one of solution hmmm it can be 4 strange....

Nnesha (nnesha):

i'll go with -3.83

Nnesha (nnesha):

good luck! check ur question one more time :-)

OpenStudy (anonymous):

Okay thanks